OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

find the slope

OpenStudy (anonymous):

y=(1/4)x+(3/2)

OpenStudy (anonymous):

m=1/4 y-3=(1/4)(x-6) 4y-12=x-6 4y-x=6

OpenStudy (anonymous):

hi again erik- with these you just find the slope, put it into the point slope form from before (just choose either of the points) and then simplify into y=mx+b form

OpenStudy (anonymous):

okay thnx for the help